# Mitigating the Accuracy-Robustness Trade-off via Multi-Teacher Adversarial Distillation

### v1

@@ 4.5 Further Explorations for Different Teacher Models
To better understand the impact of different teachers, we evaluate the performance of the same students guided by different teachers. We choose ResNet-18 as the student, and ResNet-56 and WideResNet-34-10-C (95.83% clean accuracy) as the clean teacher, respectively. Also, we choose WideResNet-34-10 and WideResNet-70-16 (From [11] with stronger robustness) as the robust teacher, and other experimental settings are the same as the initial settings. The performance is shown in Table VIII.
TABLE VIII: Study with different teacher models using our MTARD, WRN-34-10-C is the clean teacher model WideResNet-34-10-C, WRN-34-10 and WRN-70-16 are the abbreviations of WideResNet-34-10 and WideResNet-70-16. RN-56 is the abbreviation of ResNet-56. All the results are the best checkpoints based on W-Robust Acc.
Attack
Clean Teacher
Robust Teacher
W-Robust
[…]
RN-56
WRN-34-10
69.94%
WRN-34-10-C
WRN-34-10
69.09%
WRN-34-10-C
WRN-70-16
68.49%
${}_{\scriptscriptstyle\rm{\infty}}$[3]
RN-56
WRN-34-10
69.04%
WRN-34-10-C
WRN-34-10
68.48%
WRN-34-10-C
WRN-70-16
67.89%
@@ 5 Conclusion
This paper focused on mitigating the trade-off between accuracy and robustness in adversarial training. To bring both clean and robust knowledge, we proposed Multi-Teacher Adversarial Robustness Distillation (MTARD) to guide the student model, in which a clean teacher and a robust teacher are applied during the adversarial training process. In addition, we proposed the Entropy-Based Balance algorithm to keep the knowledge scale of these teachers consistent. In order to ensure that the student gets equal knowledge from two teachers, we designed a method to use Normalization Loss Balance in MTARD. A series of experiments demonstrated that our MTARD outperformed the existing adversarial training methods and adversarial robustness distillation on CIFAR-10, CIFAR-100, and Tiny-ImageNet. In the future, MTARD can be applied to other tasks with multiple optimization goals, not only limited to adversarial robustness distillation, which has great development potential.

### v3

@@ 4.6 Further Explorations for Different Teacher Models
To better understand the impact of different teachers, we evaluate the performance of the same students guided by different teachers. We choose ResNet-18 as the student, and ResNet-56 and WideResNet-34-10-C (95.83% clean accuracy) as the clean teacher, respectively. Also, we choose WideResNet-34-10 and WideResNet-70-16 (From [15] with stronger robustness) as the robust teacher, and other experimental settings are the same as the initial settings. The performance is shown in Table XI.
TABLE XI: Study with different teacher models using our B-MTARD, WRN-34-10-C is the clean teacher model WideResNet-34-10-C, WRN-34-10 and WRN-70-16 are the abbreviations of WideResNet-34-10 and WideResNet-70-16. RN-56 is the abbreviation of ResNet-56. All the results are the best checkpoints based on W-Robust Acc.
Attack Clean Teacher Robust Teacher W-Robust
PGDsat[31] RN-56 WRN-34-10 69.94%
WRN-34-10-C WRN-34-10 69.09%
WRN-34-10-C WRN-70-16 68.49%
PGDtrades[55] RN-56 WRN-34-10 71.30%
WRN-34-10-C WRN-34-10 70.66%
WRN-34-10-C WRN-70-16 70.01%
CW[4] RN-56 WRN-34-10 69.04%
WRN-34-10-C WRN-34-10 68.48%
WRN-34-10-C WRN-70-16 67.89%
@@ 5 Conclusion
This paper focused on mitigating the trade-off between accuracy and robustness in Adversarial Training. To bring both clean and robust knowledge, we proposed Balanced Multi-Teacher Adversarial Robustness Distillation (B-MTARD) to guide the student model, in which a clean teacher and a robust teacher are applied in a balanced state during the Adversarial Training process. In addition, we proposed the Entropy-Based Balance algorithm to keep the knowledge scale of these teachers consistent. In order to ensure that the student gets equal knowledge from two teachers, we designed a method to use Normalization Loss Balance in B-MTARD. A series of experiments demonstrated that our B-MTARD outperformed the existing Adversarial Training methods and adversarial robustness distillation on CIFAR-10, CIFAR-100, and Tiny-ImageNet. In the future, B-MTARD can be applied to other tasks with multiple optimization goals, not only limited to adversarial training, which has great development potential.
Acknowledgement
This work was supported by the Project of the National Natural Science Foundation of China (No.62076018), and the Fundamental Research Funds for the Central Universities.
Appendix A The Proof
Proof 1
Based on the definition of relation entropy, the knowledge scale $K^{T}$ can further expend as follows:
$K^{T}=KL(I(x),T(x))=-\sum_{k=1}^{C}p_{k}^{T}(x)log(p_{k}^{I}(x))-H(P^{T}),$ (23)
where $P^{I}=\{p_{1}^{I}(x),...,p_{C}^{I}(x)\}$ and $P^{T}=\{p_{1}^{T}(x),...,p_{C}^{T}(x)\}$ denote the general model $I$ predicted distribution and the well-trained teacher model $T$ predicted distribution, respectively. For the general model, the model’s prediction has no preference towards any samples, so a reasonable assumption can be made as follows:
$p_{k}^{I}(x)=1/C,k=1,2,...,C.$ (24)
Based on this assumption, we can easily find that the first term ($-\sum_{k=1}^{C}p_{k}^{T}(x)log(p_{k}^{I}(x))$) in Eq.(23) satisfies:
$-\sum_{k=1}^{C}p_{k}^{T}(x)log(p_{k}^{I}(x))=-\sum_{k=1}^{C}p_{k}^{T}(x)log(1/C),$ (25)
because the sum of the model’s prediction is equal to 1, so $\sum_{k=1}^{C}p_{k}^{T}(x)=1$, then we have:
$-\sum_{k=1}^{C}p_{k}^{T}(x)log(p_{k}^{I}(x))=-log(1/C)=logC,$ (26)
and $-log(1/C)$ is a constant that does not change, so the teacher’s knowledge scale is negatively related to the second term in Eq.(23) (information entropy $H(P^{T})$) as follows:
$K^{T}=logC-H(P^{T}).$ (27)
Then the Theorem 1 is proved.
Proof 2
Here we firstly define the $q$ as $exp(z(x)/\tau)$ in Eq. (3) of main text, and $p_{m}$= $q_{m}/\sum_{j=1}^{C}q_{j}$, and the distribution $P=\{p_{1},p_{2},...,p_{C}\}$ , then we further expand information entropy $H(P)$ as follows:
$H(P)=-\sum_{j=1}^{C}p_{j}log(p_{j}),$ (28)
$=-\sum_{j=1}^{C}(\frac{q_{j}}{\sum_{k=1}^{C}q_{k}})log(\frac{q_{j}}{\sum_{k=1}^{C}q_{k}}),$ (29)
$=-\frac{[(\sum_{j=1}^{C}q_{j}logq_{j})-(\sum_{j=1}^{C}q_{j})log(\sum_{j=1}^{C}q_{j})]}{\sum_{j=1}^{C}q_{j}},$ (30)
$=-\frac{(\sum_{j=1}^{C}q_{j}logq_{j})}{\sum_{j=1}^{C}q_{j}}+log\sum_{j=1}^{C}q_{j}.$ (31)
After the information entropy $H(P)$ is decomposed, we first try to solve the partial derivative of information entropy $H(P)$ with respect to $q_{m}$ as follows:
$\nabla_{q_{m}}H(P)$
$=\frac{1}{\sum_{j=1}^{C}q_{j}}-\frac{((logq_{m}+1)\sum_{j=1}^{C}q_{j}-\sum_{j=1}^{C}q_{j}logq_{j})}{(\sum_{j=1}^{C}q_{j})^{2}},$ (32)
$=-\frac{(logq_{m}\sum_{j=1}^{C}q_{j}-\sum_{j=1}^{C}q_{j}logq_{j})}{(\sum_{j=1}^{C}q_{j})^{2}},$ (33)
$=-\frac{logq_{m}}{\sum_{j=1}^{C}q_{j}}+\frac{\sum_{j=1}^{C}q_{j}logq_{j}}{(\sum_{j=1}^{C}q_{j})^{2}}.$ (34)
Then we compute the partial derivative of $q_{m}$ with respect to $\tau$ as follows:
$\nabla_{\tau}q_{m}=-\frac{q_{m}z_{m}(x)}{\tau^{2}}=-\frac{q_{m}logq_{m}}{\tau}.$ (35)
According to the chain rule of derivation, we can compute the partial derivative of information entropy with respect to $\tau$ as follows:
$\nabla_{\tau}H(P)=\sum_{m=1}^{C}\nabla_{q_{m}}H(P)\nabla_{\tau}q_{m}$ (36)
$=\sum_{m=1}^{C}[\frac{-logq_{m}}{\sum_{j=1}^{C}q_{j}}+\frac{\sum_{j=1}^{C}q_{j}logq_{j}}{(\sum_{j=1}^{C}q_{j})^{2}}](-\frac{q_{m}logq_{m}}{\tau})$ (37)
$=\frac{1}{\tau}\sum_{m=1}^{C}[\frac{q_{m}log^{2}q_{m}}{\sum_{j=1}^{C}q_{j}}-\frac{q_{m}logq_{m}\sum_{j=1}^{C}q_{j}logq_{j}}{(\sum_{j=1}^{C}q_{j})^{2}}]$ (38)
$=\frac{[(\sum_{j=1}^{C}q_{j})(\sum_{j=1}^{C}q_{j}log^{2}q_{j})-(\sum_{j=1}^{C}q_{j}logq_{j})^{2}]}{\tau(\sum_{j=1}^{C}q_{j})^{2}}$ (39)
Proof 3
Here we try to prove the $\nabla_{\tau}H(P)\geq 0$, according to Cauchy-Buniakowsky-Schwarz inequality, the derivation is as follows:
$(\sum_{j=1}^{C}q_{j})(\sum_{j=1}^{C}q_{j}log^{2}q_{j}),$ (40)
$=(\sum_{j=1}^{C}(\sqrt{q_{j}})^{2})(\sum_{j=1}^{C}((\sqrt{q_{j}})logq_{j})^{2}),$ (41)
$\geq(\sum_{j=1}^{C}q_{j}logq_{j})^{2}.$ (42)
The equal sign $=$ holds only if all $q_{j}$ are equal, and the $\tau$ and $q_{i}$ is always more than 0,
$\nabla_{\tau}H(P)$
$=\frac{(\sum_{j=1}^{C}q_{j})(\sum_{j=1}^{C}q_{j}log^{2}q_{j})-(\sum_{j=1}^{C}q_{j}logq_{j})^{2}}{\tau(\sum_{j=1}^{C}q_{j})^{2}}\geq 0.$ (43)
Then conclusion $\nabla_{\tau}H(P)\geq 0$ is proved.